OpenStudy (anonymous):

Am I correct when a degree of a whole number like (7) 0? But when you're using it in an exponent property for example, 7x7^5. Why would the answer be 7^6? as in the first 7 is 7^1. Is the degree of a number and an exponent two different things? I feel like this is too confusing...... :P

OpenStudy (anonymous):

of course it is confusing because it means two different things. the degree of a polynomial is the highest power of a variable, not a number

OpenStudy (anonymous):

so \{x^2+3x+64\) has degree 2, whereas 64 has degree zero (since it is \(64x^0\) ) and not for example degree 6 because \(64=2^6\) or degree 3 because \(64=4^3\) or degree 2 because \(64=8^2\) or degree 1 because \(64=64^1\)

OpenStudy (anonymous):

first line should read \(x^2+3x+64\) has degree 2

OpenStudy (anonymous):

Oh alright, that makes sense. Degrees pertain to how many variables? as in x^2 is the highest in that polynomial because it's two x's. and 64 is zero because it has no x's right?